OpenStudy (anonymous):

so i got a question when it ask for the diameter of a circle and it gives me the radius do i put for example 7pi or just 7 and im confused with the whole circle/diamerter thing can someone explain please

OpenStudy (butterflydreamer):

Hint: Diameter = 2 * radius

OpenStudy (butterflydreamer):



OpenStudy (anonymous):

uhh im still lost for example like diameter is 7 wht is the circumference of a circle wht do i put

OpenStudy (butterflydreamer):

ahhhh okay! So do you know what the formula is to find the circumference of a circle? :)

OpenStudy (anonymous):

kinda

OpenStudy (butterflydreamer):

\[circumference= 2 \pi r\] Where r = radius. Remember how i said diameter = 2 * radius? this means that d= 2r Therefore you want to find the value of "r" first :) r = 1/2 * d

OpenStudy (anonymous):

like tht for example Suppose the circumference of a circle is 16π. What is its radius?

OpenStudy (anonymous):

this i mean this

OpenStudy (butterflydreamer):

\[Circumference = 2 \pi r \] They've told us that the circumference of the circle equals to 16 pi and we want to find the radius right? \[Circumference = 16 \pi \] So this means that \[2 \pi r = 16 \pi\] Now just solve for "r" (radius)

OpenStudy (anonymous):

soooo 6pi

OpenStudy (anonymous):

o no i mean 8pi

OpenStudy (butterflydreamer):

close but not quite! \[2 \pi r = 16 \pi\] Divide both sides by 2 pi \[\frac{ 2 \pi r }{ 2 \pi} = \frac{ 16 \pi }{ 2 \pi}\] \[r = 8 \]

OpenStudy (anonymous):

o ok soo i got anotherone Suppose the circumference of a circle is 3π. What is its diameter?

OpenStudy (butterflydreamer):

follow the same pattern. Circumference of a circle formula: \[Circumference = 2 \pi r\] We are given the following information: \[circumference = 3 \pi\] \[\therefore 2 \pi r = 3 \pi\] Divide both sides by 2 pi to find "r". After you find "r", remember that d = 2r. Since we want to find diameter, multiply r by 2.

OpenStudy (anonymous):

3?

OpenStudy (butterflydreamer):

excellent!! :D

OpenStudy (anonymous):

awesome thx and another one a diameter is 12 what is the radius

OpenStudy (butterflydreamer):

no worries :) remember this: DIAMETER = 2 TIMES RADIUS. RADIUS = 1/2 DIAMETER So if d = 2r And we're told d = 12 12 = 2r Solve for r